3^1+x+4.3^-3-7=0
1
Expert's answer
2013-07-10T08:03:05-0400
3^1+x+4.3^-3-7=0
3+x+(1/(4.3)^3)-7=0

whence x=4-(1/(4.3)^3)=approximately 3.99